OpenStudy (anonymous):

Simplify the following expression. 6x2 + 14x + 8 3x + 4

OpenStudy (anonymous):

6x^2+14x+8 =2(3x^2+7x+4) =2(3x+4)(x+1) Through trinomial factoring